MODS1 Commissioning Run 7: UTC 2011 April 13

Observer(s): R. Pogge, R. Stoll (OSU), E. Skillman (UMinn), M. Fumana (INAF), M. Wagner (OSU/LBTO)
Telescope Operator: G. Bechetti (LBTO)
AO Support: Remote from Tucson as needed
Telescope Support: J. Little (LBTO)
Instrument Support: D. Officer (LBTO)

Summary

The LDG Baffle works great, we are now in position to realize the true potential of MODS1. Observed all night in moderately good (1-1.5arcsec) seeing but with some wind. MODS1 came back to life immediately after the baffle installation and worked all night without problems.

Preparations

The primary task during the day was to install the new Left Direct Gregorian Baffle. This required us to dismount MODS1, leave it dangling in mid-air on the crane (yeah, really, we have pictures), install the baffle in the Left Primary Mirror Cell, then re-mount MODS1. The process took from about 8:30am until 1:30pm, including a practice installation of the baffle on the Right Primary Mirror Cell before we committed to dismounting MODS1. Help from John Little, Rick Hanson, and Jarrod Schuh made the job go great.

Only one hitch on dismount: the glycol return fitting (male) labeled R1 on the LDG rotator is leaking from the inside seal when demated from the MODS1 line. This spilled a bit of glycol on the instrument, the floor, and me. Not so self-sealing anymore. IssueTrak #3314 was filed with the report. We can work with it, but it should be replaced: remember to turn off the glycol supply and return to those feeds and carry absorbent towels when the time comes to avoid spilling more glycol onto MODS1 (and yourself).

After the remount, we reconnected MODS1 and it came up without problems. IMCS was fired up in Zenith-pointing dual-imaging configuration and was spot on alignment: whatever there is about the rotator that distorts MODS from the on-cart configuration appears to be repeatable. We will gather more detailed information this week.

A first shot using full dome lights and the hatch open with a full-frame read on both detectors with the clear filter and the dark mask blocking the direct view to the collimator (worst-case before for seeing stray light getting direct into the cameras from scattering inside the mirror cell) shows little remaining residual structure only at the few ADU level, quite unlike the many 1000s of ADU and rich structure we were seeing before. A detailed comparison will be undertaken as part of this week's work.

Details

All times are UTC unless otherwise noted.

Tonight was a lot of instruction and explanation while exercising the instrument after the baffle installation, so I have not been keeping a chronological wiki as before. I will attempt to reconstruct from hand-written notes and the image logs.

The main goal was to heavily exercise target acquisition scripts for long-slit spectra, getting the other observers to write them and execute them for various targets ranging from routine (bright spectrophotometric standard stars) to challenging (an r=22 faint blue galaxy which is dominated by emission lines). With the low background afforded by the baffle despite the moon, we also pushed some fainter targets to demonstrate that the baffles worked. And they do, beautifully. We will examine the many spectra from tonight to quantify the background levels and their greater uniformity.

The bottom line on the baffle is that all of the bad scattered light we've seen previously from bright moonlight - undispersed parasitic light in spectra clobbering faint spectra, non-uniform background in spectra and images, glints and bright spots in the acquisition/guide camera have been eliminated. As the moon gets brighter through the week we will stress the system further. Since this is a primarily dark-time instrument, we are now at a place where we can start addressing deferred issues like sky background, sky flats, slit illumination correction images, scattered light between multi-object slitlets, etc.

A second goal was to collect as many spectrophotometric stars in grating and prism mode to start building the exposure time matrix for the standard stars most observers will use for grating and prism modes. These will form the library of standard stars for partner observing.

A few brief observations:

Getting going at the start of the night was very slow, with John Hill assisting from Tucson. Us instrument folks had nothing to do until I was asked to reboot the AzCamServer for the guide camera, apparently it came up in a wedged state. This seemed to take a great deal of time and it was not until 0300UTC that we acquired our first pointing star.

For much of the night active optics was converging very slowly. After we came near mirror limits and had to unwrap the azimuth, things behaved much better for an hour or so after Geno cleared active optics. This happened a couple of other times during the night. Clearing active optics and restoring basic function seemed to take a rather long time to complete, mostly because the telescope operator is less facile at this aspect of operating the LBT.

Wind was a bother but never came close to shutting us down. Conditions were mostly clear with periods of thin cirrus blowing over. Mostly it meant we had problems of dome seeing because we could not open vent doors and flush warm air out of the enclosure.

Seeing was 1.5 arcsec with brief periods of 0.8 arcsec seeing.

-- RichardPogge - 12 Apr 2011
Topic revision: r3 - 13 Apr 2011, RichardPogge
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback